John has 100 pages to read over 5 days. He wants to increase the number of pages he reads by 5 pages each day.
a. How many pages should he read the first day?
b. John learns that he has 200 pages to read. If he reads twice as many pages the first day, then increases the number of pages he reads by 5 pages a day, will he read twice as many pages? Explain your reasoning.

Answer by ewatrrr(24785)

a. x + (x+5) + (x+10) + (x+15) + (x +20) = 100 pages
5x + 50 = 100
x = 10
b. IF x = 20
5x+ 50 = 150. NO, will NOT read twice as many pages